OpenStudy (anonymous):

@Mertsj Can you help me set up a 5 step problem -thing?

OpenStudy (mertsj):

I don't know what a 5 step problem thing is.

OpenStudy (anonymous):

Let me copy and paste it. :P

OpenStudy (anonymous):

The first 3 steps say this: Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: Drive a boat Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. I put: Monica is driving a boat at 40 mph, for 2 hours. But, what confuses me is the way they word Step 4. Step 4 says: Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). And here is step 5: Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (mertsj):

Have to wait...I'm on the phone

OpenStudy (anonymous):

Alright. Will you be very long? I can ask someone else if you're too busy.

OpenStudy (anonymous):

@ash2326 could you help?

OpenStudy (kymber):

Step four is asking how long it takes them when they're not going against the current - like the current in the ocean, the direction the water is moving

OpenStudy (anonymous):

Would I have to make that up?

OpenStudy (kymber):

Yeah, I think it just wants you to pick a smaller amount of time

OpenStudy (anonymous):

So, I could say Monica is driving a boat at 40 mph, for 2 hours. And Mike was doing the same, but only for 1 hour?

OpenStudy (anonymous):

Whats the current speed?

OpenStudy (mertsj):

Are you done with me?

OpenStudy (anonymous):

lets assume the current speed is 8mph. the boat is driven at 40 mph against the current you assumed he was driving for 2 hours the trip (d) is calculated like this: s = d/t => (40-8) = d / 2 => 3 = 2*(40-8) d= 64 miles for the same trip (64miles) now with the flow you want to know the time. s = d/t => (40+8) = 64/t => t=64/(40+8) t= 1.33 hours ~ 1h20h for standing still water and for the same trip (64miles) t = 64/40 = 1.6h ~ 1h36min

OpenStudy (anonymous):

Not really @Mertsj I'm still confused.